OpenStudy (anonymous):

HELP!!!!!! Solve. x2 + 5x – 6 = 0 A. {1, –6} B. {2, 3} C. {–1, 6} D. {–2, –3}

OpenStudy (anonymous):

\[x ^{2}+5x-6 =0\] is what you mean right?

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

oh sorry. no its A

OpenStudy (anonymous):

Yeah A is correct :D

OpenStudy (anonymous):

for this, you have a variable with a squared to start off with! this means you are going to have two answers. if i said x^2 equals 9 you automatically think 3 is the answer. -3 can also be though! to solve an equation like this where it is equal to zero, you wanna set it up in this form! (x )(x )=0

OpenStudy (anonymous):

okay thank you and @jeremyggg can you keep explaining it?

OpenStudy (anonymous):

you wanna think of your last number of -6. what two numbers can you multiple by that will equal -6 but when added together will equal 5?

OpenStudy (anonymous):

+6 and -1 right? 6*-1 = -6 and 6+-1 = 5 so those are your two numbers you add to the (x )(x ) =0 (x + 6)(x-1)=0

OpenStudy (anonymous):

from here, we know that any number *0=0 soooo there are two numbers that you can use to make either of those equal 0! if x was 1 the (x-1) would be 0! or.... if x was -6 (x+6) would equal 0!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

seems tough, but the steps for anything x^2 +x + c = 0 1)put it in this form (x )(x ) = 0 2)then make each parenthesis equal 0!